Madera's mayor is watching a car from the ledge of the water tower. The car is directly approaching the water tower. When first noticed, the angle of depression to the car is 58°. When the car stops, the angle of depression is 77°. The mayor's line of sight is 127 feet above the ground.


Part 1

How far did the car travel from when it was first noticed until it stopped? Round to the nearest foot. 

The car traveled ___________ feet from when it was first noticed until it stopped.


Part 2

Explain how you determined the distance the car had traveled by applying your understanding of the angles and the height of the mayor standing on top of the water tower. 






1
Expert's answer
2021-12-16T15:30:18-0500

Solution:


I) We should find the distance at each given angle:


1) "\frac{127}{tan(58\degree)}=79.36 ft"


2) "\frac{127}{tan(77\degree)}= 29.32ft"


In order to find the distance car travel from when it was first noticed until it stopped:


79.36 - 29.32 = 50.04 feet

Answer: 50 feet



II) The type of math it's using is trigonometric ratios and Pythagoras theorem.